# Modeling User Intent Beyond Trigger: Incorporating Uncertainty for Trigger-Induced Recommendation

### v1

@@ Abstract.
Click-Through Rate Prediction; Trigger-Induced Recommendation; Contrastive Learning; User Intent Modeling
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3680065isbn: 979-8-4007-0436-9/24/10ccs: Information systems Recommender systemsccs: Information systems Personalizationccs: Information systems Learning to rank
@@ 1. Introduction
To address these challenges, we propose a novel model in this paper called DUIN  [^2] The source code of DUIN is released at: https://github.com/majx1997/DUIN., short for Deep Uncertainty Intent Network. It primarily comprises three components. To extract explicit user intent based on the trigger, we design an Explicit Intent Exploit Module (EIEM) that uses the contrastive learning paradigm to obtain generalizable and distinguishable intent representations. We also design a Latent Intent Explore Module (LIEM) to effectively explore the users’ underlying intent by leveraging the multi-view relationships between items, applicable to both the trigger item and the target item. An Intent Uncertainty Measurement Module (IUMM) is implemented to model user intent intensity as a Gaussian distribution, capturing the uncertainty associated with user intent. Overall, the contributions of this paper can be summarized as follows:
@@ 3.2.1. Explicit Intent Exploit Module
User chronological behaviors exhibit causality and convey explicit intent information, including aspects users are inclined to interact with, such as brand, theme, and category. The trigger item merely represents a partial expression of user explicit intent. In DUIN, we introduce the Explicit Intent Exploit Module to learn user explicit intent.
@@ 3.3. Prediction and Optimization
Having obtained the user intent representations $\mathcal{H}_{ei}$, $\mathcal{H}_{i}$ and $\mathcal{H}_{li}$ from both EIEM and LIEM, similar to equation 1, we concatenate these intent representations with other features, including context features $u^{c}$, profile features $u^{p}$, the target item $i^{tr}$, and others. Subsequently, the concatenated features are fed into MLP (Zhou et al., 2018) to estimate the likelihood $\hat{y_{i}}\in[0,1]$ of user interaction. We adopt the binary cross entropy loss (Zhou et al., 2018; Shen et al., 2022; Wang et al., 2017) for model optimization:
@@ 4.1. Experimental Setup
In offline experiments, we use commonly used metrics AUC and RelaImpr (Zhou et al., 2019; Xiao et al., 2020; Feng et al., 2019), which are widely recognized evaluation metrics in CTR tasks. We conduct each experiment five times and report the average results. All other experimental and parameter settings remain consistent with  (Shen et al., 2022; Xiao et al., 2024) to ensure fair comparison. One-sided Wilcoxon rank-sum p-value is calculated to quantify the significance of the improvement achieved by DUIN (Gehan, 1965). In online A/B testing, we use CTR and Conversion Rate (CVR) as online metrics to assess the efficacy of DUIN.
@@ 4.3.2. Online A/B Testing Results
As shown in Table 4, we conduct online A/B testing to evaluate DUIN in Detail Recommendation and Mini Detail, denoted as TIR1 and TIR2 respectively. The Daily Active Users (DAU) for the two TIR scenarios are approximately 700K and 300K, respectively. We choose DEI2N (Xiao et al., 2024) as the baseline, as it performs the second-best in offline tests. We conduct 14-day experiments in the two scenarios, and the experimental results are statistically significant according to an unpaired t-test. Results show that DUIN consistently outperforms DEI2N in terms of CTR and CVR. Furthermore, since the exposed items can be classified into two categories—those that belong to the same category as the trigger item and those that belong to a different category, CTR can also be segmented into Same-category CTR (S-CTR) and Cross-category CTR (C-CTR). DUIN also shows significant improvements under both S-CTR and C-CTR metrics, indicating its superior intent representation capabilities. From the perspective of cross-category exposure ratio (CCER), DUIN alleviates the problem of over-convergence in intention understanding and provide users with more diverse and novel recommendations. Due to DUIN’s excellent online performance, we have deployed it in major TIR scenarios on the Alibaba.com platform.
@@ 5. Conclusion
In this paper, we introduce a novel TIR model, called Deep Uncertainty Intent Network (DUIN), to model both explicit and latent user intent with uncertainty based on user-triggered items. The proposed DUIN consists of three modules: i) Explicit Intent Exploit Module, which aims to obtain generalizable and distinguishable intent representations; ii) Latent Intent Explore Module, designed to explore the user’s underlying intent; and iii) Intent Uncertainty Measurement Module, tasked with modeling user intent intensity as a distribution to capture the uncertainty associated with user intent. Extensive experiments demonstrate that DUIN surpasses various representative methods, achieving state-of-the-art performance. Moreover, DUIN has been deployed across all TIR scenarios within our Alibaba.com platform, and online A/B testing has conclusively verified the superiority of our approach.

### v2

@@ Abstract.
Click-Through Rate Prediction, Trigger-Induced Recommendation, Contrastive Learning, User Intent Modeling
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Personalizationccs: Information systems Recommender systemsccs: Information systems Learning to rank
@@ 1. Introduction
To address these challenges, we propose a novel model in this paper called DUIN  [^2] The source code of DUIN will be released at: https://github.com/majx1997/DUIN., short for Deep Uncertainty Intent Network. It primarily comprises three components. To extract explicit user intent based on the trigger, we design an Explicit Intent Exploit Module (EIEM) that uses the contrastive learning paradigm to obtain generalizable and distinguishable intent representations. We also design a Latent Intent Explore Module (LIEM) to effectively explore the users’ underlying intent by leveraging the multi-view relationships between items, applicable to both the trigger item and the target item. An Intent Uncertainty Measurement Module (IUMM) is implemented to model user intent intensity as a Gaussian distribution, capturing the uncertainty associated with user intent. Overall, the contributions of this paper can be summarized as follows:
@@ 3.2.1. Explicit Intent Exploit Module
User chronological behaviors exhibit causality and convey explicit intent information, including aspects users are inclined to interact with, such as brand, theme and category. The trigger item merely represents a partial expression of user explicit intent. In DUIN, we introduce the Explicit Intent Exploit Module to learn user explicit intent.
@@ 3.3. Prediction and Optimization
Having obtained the user intent representations $\mathcal{H}_{ei}$, $\mathcal{H}_{i}$ and $\mathcal{H}_{li}$ from both EIEM and LIEM, similar to equation 1, we concatenate these intent representations with other features, including context features $u^{c}$, profile features $u^{p}$, and the target item $i^{ta}$. Subsequently, the concatenated features are fed into MLP (Zhou et al., 2018) to estimate the likelihood $\hat{y_{i}}\in[0,1]$ of user interaction. We adopt the binary cross entropy loss (Zhou et al., 2018; Shen et al., 2022; Wang et al., 2017) for model optimization:
@@ 4.1. Experimental Setup
In offline experiments, we use commonly used metrics AUC and RelaImpr (Zhou et al., 2019; Xiao et al., 2020; Feng et al., 2019), which are widely recognized evaluation metrics in CTR tasks. We conduct each experiment five times and report the average results. All other experimental and parameter settings remain consistent with  (Shen et al., 2022; Xiao et al., 2024) to ensure fair comparison. One-sided Wilcoxon rank-sum p-value is calculated to quantify the significance of the improvement achieved by DUIN (Gehan, 1965). In online A/B testing, we use CTR and Conversion Rate (CVR) as online metrics to assess the efficacy of DUIN. All reproducible materials of DUIN will be available after paper review.
@@ 4.3. Online A/B Testing
We have implemented our proposed method, DUIN, in industrial Trigger-Induced Recommendation scenarios to serve substantial traffic requests. In this section, we will describe the online A/B testing settings and results.
@@ 4.3.2. Online A/B Testing Results
We conduct online A/B testing to evaluate DUIN in two online TIR scenarios over a 14-day period, as shown in Table 4. We choose DEI2N (Xiao et al., 2024) as the baseline, which performs the second-best in offline tests. Results show that DUIN consistently outperforms DEI2N in terms of CTR and Conversion Rate (CVR). Furthermore, since the exposed items can be classified into two categories—those that belong to the same category as the trigger item and those that belong to a different category, CTR can also be segmented into Same-category CTR (S-CTR) and Cross-category CTR (C-CTR). DUIN shows significant improvements under both S-CTR and C-CTR metrics, which also shows that DUIN has better intent representation capabilities. From the perspective of cross-category exposure ratio (CCER), DUIN can alleviate the problem of over-convergence of intention understanding states and provide users with more diverse and novel recommendations. Due to DUIN’s excellent online performance, we have deployed it in major TIR scenarios on Alibaba.com platform.
@@ 5. Conclusion
In this paper, we introduce a novel TIR model, called Deep Uncertainty Intent Network (DUIN), to model both explicit and latent user intent with uncertainty based on user-triggered items. The proposed DUIN model consists of three modules: i) Explicit Intent Exploit Module, which aims to obtain generalizable and distinguishable intent representations; ii) Latent Intent Explore Module, designed to explore the user’s underlying intent; and iii) Intent Uncertainty Measurement Module, tasked with modeling user intent intensity as a distribution to capture the uncertainty associated with user intent. Extensive experiments demonstrate that DUIN surpasses various representative methods, achieving state-of-the-art performance. Moreover, DUIN has been deployed across all TIR scenarios within our Alibaba.com platform, and online A/B testing has conclusively verified the superiority of our approach.